OpenStudy (jody.wilson):

If a polynomial has three terms, x2 + 12x + 36, which factoring method can be considered?

OpenStudy (hyuna301):

usually there are 3 methods but since the equation is easy i will use the factorisation method..

OpenStudy (hyuna301):

given eq. , \[x ^{2} + 12 x + 36 = 0\] \[x ^{2} + 6x + 6x + 36 = 0\] \[( x + 6 ) ( x + 6) = 0\]